two forces act on an object. a 36 n force acts at 225° , A 48-N force acts at 315°. what would be tge magnitude and directionof their equilibrant?


1
Expert's answer
2021-09-14T18:57:52-0400

Let's first find "x"- and "y"-components of resultant force:


"F_x=36\ N\cdot cos225^{\circ}+48\ N\cdot cos315^{\circ}=8.48\ N,""F_y=36\ N\cdot sin225^{\circ}+48\ N\cdot sin315^{\circ}=-59.4\ N."


We can find the magnitude of the resultant force acting on an object from the Pythagorean theorem:


"F=\sqrt{F_x^2+F_y^2}=\sqrt{(8.48\ N)^2+(-59.4\ N)^2}=60\ N."

We can find the direction of the resultant force acting on an object from the geometry:


"\theta=sin^{-1}(\dfrac{F_y}{F})=sin^{-1}(\dfrac{-59.4\ N}{60\ N})=278^{\circ}."


The resultant force directed at "278^{\circ}" counterclockwise from the positive "x"-axis.